OpenStudy (anonymous):

The mean of the data set is 11. What is the value of ?

OpenStudy (anonymous):

question is incomplete :)

OpenStudy (anonymous):

it is a weird formula with a roman E

OpenStudy (anonymous):

Sum notation. You mean sigma?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

You need to type the rest of the question

OpenStudy (anonymous):

6 - ∑(Xi - x) i=1

OpenStudy (anonymous):

find the deviation from the mean of each Xi, then add them all up

OpenStudy (anonymous):

say X1 = 6, then X1 - 6 = 6-6 = 0 X2 = 10, then X2 - 6 = 10-4 = 4 0 + 4 + (X3 - 6) +... + (X6 - 6)